Moon (Ltd) manufacture specially treated garden benches. The following information was extracted from the budget for the year ended 29 February 2016: Estimated sales for the financial year 2 000 units Selling price per garden bench R450 Variable production cost per garden bench: - Direct material - Direct labour - Overheads R135 R90 R45 Fixed production overheads R127 500 Selling and administrative expenses: - Salary of sales manager for the year - Sales commission R75 000 10% of sales Required: (round off answers to the nearest rand or whole number) 3.1 Calculate the break-even quantity. (4 marks) 3.2 Determine the break-even value using the marginal income ratio. (4 marks) 3.3 Calculate the margin of safety (in Rand terms). (4 marks) 3.4 Determine the number of sales units required to make a profit of R150 000. (3 marks) 3.5 Suppose Moon (Ltd) wants to make provision for a 10% increase in fixed production costs and an increase in variable overhead costs of R15 per unit. Calculate the new break-even quantity. (5 marks)
Answer #1

3.1) Break Even quantity = Fixed cost/Contribution Margin Per unit

=202,500/135

=1,500 Units

3.2)break-even value = Fixed cost/Marginal Income ratio

=202,500/30%

=675,000

3.3)Margin of Safety = 900,000 - 675,000

= 225,000

3.4)Sales unit Required to earn 150,000 = (150,000+202,500)/135

=2611.11 units

3.5)New Contribution = 135 -15 = 120

New fixed cost = 202,500 + (127,500 x 10%) = 215,250

New Breakeven quantity = 215,250/120 = 1,793.75

Working note

Sales $          450.00
Variable cost $          315.00
Contribution Margin $          135.00
Fixed cost $ 202,500.00